A total of $35,000 is invested into two banks accounts paying 6% and 4% simple interest. The yearly interest is $1710. How much is invested at each rate?
Answer by mananth(16946)

Part I 6.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 35000 -------------1
Interest on both investments
6.00% x + 4.00% y= 1710
Multiply by 100
6 x + 4 y= 171000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -210000.00
Add this to (2)
0 x -2 y= -39000
divide by -2
y = 19500
Part I 6.00% $ 15500
Part II 4.00% $ 19500

CHECK
15500 --------- 6.00% ------- 930.00
19500 ------------- 4.00% ------- 780.00
Total -------------------- 1710.00

m.ananth@hotmail.ca